# Empowering Agile-Based Generative Software Development through Human-AI Teamwork

### v1

@@ 7. Related Work
This paper first introduces an AI agent for generative software development (AgileGen) based on an Agile methodologies and through human-AI Teamwork. It allows end-users and AI to handle tasks they excel at, collaboratively piecing together the development process. The related work is as follows:
@@ 7.2. Generative Software Development
Large language models such as ChatGPT (Bang et al., 2023) have been widely applied in software development and demonstrated potential. Microsoft researchers Bubeck et al. (Bubeck et al., 2023) guide GPT-4 to the development of complex HTML games. Some Python project generators based on GPT have been released on Huggingface (Generator, [n. d.]), generating function architecture based on brief requirements. AutoGPT (AutoGPT, [n. d.]) has been explored for generating HTML projects without manual prompting. Intelligent tools can help developers be more productive, but they need precise models and expert knowledge to work well. Similar projects include DemoGPT [^28] https://github.com/melih-unsal/DemoGPT, which utilizes the Streamlit visualization tool and can rapidly transform a single prompt into an AI software with the prompted functionality. This fully autonomous AI operates without human intervention in its internal processes and is prone to issues such as process freezing. Then, some questioning-based agents have been proposed. GPT-Engineer (Gpt-engineer, [n. d.]) is a project that automates the software engineering process using GPT-4, initially clarifying requirements through questions to the user before automatically generating code. Similarly, gpt-pilot [^29] https://github.com/Pythagora-io/gpt-pilot, based on GPT-4, generates software code after asking developers to clarify questions. These methods aim to assist developers; thus, the questions often require professional domain knowledge.
@@
Recently, the concept of multi-agent collaboration has been introduced. Generative Agents (Park et al., 2023) simulate human behavior by allowing agents to interact collaboratively through dialogue. CodeAgent (Tang et al., 2024) conducts code reviews through the collaboration of multiple agents led by a QA-checker. Inspired by this collaborative interaction among agents, MetaGPT (Hong et al., 2023) utilizes standard operating procedures to coordinate a multi-agent system based on large language models, with different agents simulating different software development roles. Subsequently, MetaGPT proposed a solution for multi-agent collaboration in data science applications (Hong et al., 2024). ChatDev first version (Qian et al., 2023a) designs a dialogue chain-guided, multi-agent collaborative software development agent. Based on the original prototype of ChatDev, two versions were updated. The ChatDev second version (Qian et al., 2023b) adopts experiential co-learning, collecting shortcut-oriented experiences from their historical trajectory for use in software development. The ChatDev third version (Qian et al., 2024a) refines these historical experiences through iteration before using them in software development. The fourth version of ChatDev (Qian et al., 2024b) uses a directed acyclic graph (DAG) to organize agents. Employing topological sorting simplifies agent interaction reasoning and enhances collaboration among agents. These agents follow the waterfall model, characterized by heavy documentation and process emphasis. However, they offer lower levels of human participation and control, and the hallucination errors of each agent can cascade, making it challenging to generate software products that meet user requirements. As a result, these tools are not user-friendly for end-users who wish to tailored software in software development.

### v2

@@ 7. Related Work
This paper introduces AgileGen, an AI agent designed for generative software development based on Agile methodologies and human-AI collaboration. AgileGen enables both end-users and AI to focus on the tasks they excel at, working together to effectively complete the development process. The related works are as follows:
@@ 7.2. Generative Software Development
Large language models such as ChatGPT (Bang et al., 2023) have been widely applied in software development and demonstrated potential. Microsoft researchers Bubeck et al. (Bubeck et al., 2023) guide GPT-4 to the development of complex HTML games. Some Python project generators based on GPT have been released on Huggingface (Generator, [n. d.]), generating function architecture based on brief requirements. AutoGPT (AutoGPT, [n. d.]) has been explored for generating HTML projects without manual prompting. Intelligent tools can help developers be more productive, but they need precise models and expert knowledge to work well. Similar projects include DemoGPT [^28] https://github.com/melih-unsal/DemoGPT, which utilizes the Streamlit visualization tool and can rapidly transform a single prompt into an AI software with the prompted functionality. This fully autonomous AI operates without human intervention in its internal processes and is prone to issues such as process freezing.
@@
Recently, the concept of multi-agent collaboration has been introduced. Generative Agents (Park et al., 2023) simulate human behavior by allowing agents to interact collaboratively through dialogue. CodeAgent (Tang et al., 2024) conducts code reviews through the collaboration of multiple agents led by a QA-checker. Inspired by this collaborative interaction among agents, MetaGPT (Hong et al., 2023) utilizes standard operating procedures to coordinate a multi-agent system based on large language models, with different agents simulating different software development roles. Subsequently, MetaGPT proposed a solution for multi-agent collaboration in data science applications (Hong et al., 2024). ChatDev first version (Qian et al., 2023a) designs a dialogue chain-guided, multi-agent collaborative software development agent. Based on the original prototype of ChatDev, two versions were updated. The ChatDev second version (Qian et al., 2023b) adopts experiential co-learning, collecting shortcut-oriented experiences from their historical trajectory for use in software development. The ChatDev third version (Qian et al., 2024a) refines these historical experiences through iteration before using them in software development. The fourth version of ChatDev (Qian et al., 2024b) uses a directed acyclic graph (DAG) to organize agents. Employing topological sorting simplifies agent interaction reasoning and enhances collaboration among agents. These agents follow the waterfall model, characterized by heavy documentation and process emphasis. However, they offer lower levels of human participation and control, and the hallucination errors of each agent can cascade, making it challenging to generate software products that meet user requirements.
Some questioning-based agents have been proposed to mitigate the accumulation of hallucination errors. These agents present problems that are difficult for them to solve back to humans in the way of questions, allowing humans to clarify or modify them. GPT-Engineer (Gpt-engineer, [n. d.]) is a method for automated software generation using GPT-4, which clarifies requirements with users through questions before generating code, including issues like database requirements and API usage. Similarly, gpt-pilot  [^29] https://github.com/Pythagora-io/gpt-pilot generates software code after asking developers to clarify various domain-specific questions. Data Interpreter (Hong et al., 2024) uses code to solve data science-centered analytical problems. In its dynamic plan management module, two strategies are designed: self-debugging and human editing, where tasks that fail multiple self-debugging attempts are modified through human editing. These methods aim to assist professional developers who usually require domain knowledge. End-users with business requirements are often limited by domain knowledge and find it challenging to describe requirements from a development perspective. Our approach enables end-users and agents to collaborate effectively, each handling what they excel at while jointly completing software development.